# Time-Series Forecasting and Sequence Learning Using Memristor-based Reservoir System

### v1

@@ Abstract.
Pushing the frontiers of time-series information processing in ever-growing edge devices with stringent resources has been impeded by the system’s ability to process information and learn locally on the device. Local processing and learning typically demand intensive computations and massive storage as the process involves retrieving information and tuning hundreds of parameters back in time. In this work, we developed a memristor-based echo state network accelerator that features efficient temporal data processing and in-situ online learning. The proposed design is benchmarked using various datasets involving real-world tasks, such as forecasting the load energy consumption and weather conditions. The experimental results illustrate that the hardware model experiences a marginal degradation ($\sim$4.8%) in performance as compared to the software model. This is mainly attributed to the limited precision and dynamic range of network parameters when emulated using memristor devices. The proposed system is evaluated for lifespan, robustness, and energy-delay product. It is observed that the system demonstrates a reasonable robustness for device failure below 10%, which may occur due to stuck-at faults. Furthermore, 246$\times$ reduction in energy consumption is achieved when compared to a custom CMOS digital design implemented at the same technology node.
@@ 1. Introduction
Empowering edge devices with the capability to process and learn continuously locally on the device is essential when handling stationary and non-stationary time-series data as remote cloud services has proven to be slow and unsecured (Kim et al., 2019). While such an objective enhances reliability and adaptability to edge constraints, it is challenging to satisfy in devices with limited resources. In this work, the local processing and learning from time-series data is explored within the context of reservoir networks, particularly echo state network (ESN) (Jaeger, 2001). ESN is developed to handle continuous data of stationary and non-stationary nature and has powered a large number of applications, including prosthetic finger control (Zyarah et al., 2023), stock market prediction (Wang et al., 2021), cyber attack and anamoly detection (Hamedani et al., 2017; Ullah et al., 2022), weather forecasting (De et al., 2023), and modeling dynamic motions in bio-mimic robots (Soliman et al., 2021). The network features efficient temporal data processing and unrivaled training speed (Cerina et al., 2020; Bianchi et al., 2020). Furthermore, it does not suffer from the known training problems in most recurrent neural networks, such as vanishing and exploding gradients, which occur due to the recurrent connections and gradient backward propagation through time (Zhang and Vargas, 2023). Thus, running ESN on edge devices using dedicated custom-designed hardware not only enables the algorithm to make temporal decisions in real-time but also yields a significant improvement in power efficiency and throughput (Kudithipudi et al., 2016; Gauthier et al., 2021).
@@
These aforementioned reasons urge the research community to investigate the physical implementation of ESN. There are several digital (Honda and Tamukoh, 2020; Kleyko et al., 2020; Gan et al., 2021) and mixed-signal implementations of the ESN networks in the literature using memristors devices, MOSFET switching, and photonics (Sorokina, 2020; Kume et al., 2020; Wlazlak et al., 2020; Zhang and Vargas, 2023). However, this work will place emphasis on the memristor-based implementations, owing to their numerous advantages over other approaches, such as the integration with CMOS devices and fabrication using the standard foundry process (Jiang et al., 2019). Furthermore, the memristor intrinsic non-linearity feature enhances the reservoir separability and echo state properties (Moon et al., 2019). Recent attempts of memristive ESN, including a hardware implementation of the ESN network using a memristor double-crossbar array, is presented by Hassan et al. in 2017. The proposed design performance is evaluated for autonomous signal generation using the Mackey-Glass dataset and the simulation is done in MATLAB. In 2019, Wen et al. presented a memristor-based ESN trained in online fashion using least mean square (Wen et al., 2018). The network is evaluated for short-term power load forecasting. In the same year, a memristor-based reservoir computing system, which leverages the concept of the virtual node, was introduced by Moon at al. (Moon et al., 2019). The virtual node concept involves building a single reservoir node and using it to virtually emulate a chain of other nodes in the reservoir. The proposed system features high power efficiency but suffers from high latency and short lifespan, and it is demonstrated in speech recognition and time-series forecasting. In 2022, the same concept is adopted to build a cyclic reservoir computing system, which is based on memristor devices. The proposed design is implemented in a 65nm technology node and verified on hand-written vowel recognition task (Liang et al., 2022). Lately, Nair et al. proposed a memristive ESN with an extended synaptic sampling machine (SSM) to ensure synaptic stochasticity and power efficiency (Nair et al., 2023). The proposed neural system is benchmarked for classification tasks using ECG, MNIST, Fashion MNIST, and CIFAR10.
@@
While a few of the aforementioned ESN implementations are equipped with in-situ learning to endow the ESN network with fast learning and adaptation capabilities, none extend the learning to cover the structural plasticity (synaptogenesis), which is imperative for controlling the network’s sparsity level. To the best of our knowledge, few of them utilize leaky-integrated discrete-time continuous-value neurons in the reservoir layer. Having leaky-integrated neurons is essential when dealing with temporal data as it controls the speed of the reservoir update dynamics and eventually the duration of the short-term memory of ESN (Lukoševičius, 2012). Besides the lack of using leaky-integrated neurons, none of the aforementioned implementations take into consideration the limited precision of signals digitization and undersired leakage of the sample and hold circuits during network training and testing. This paper addresses these problems with key contributions summarized as follows:
@@
The rest of the paper is organized as follows: Section II discusses the theory of ESN network and training procedure. Section III presents the system design and implementation of the ESN. The design methodology is introduced in Section IV. Section V and VI discuss the results and conclude the paper, respectively.
Figure 1. A high-level diagram of the ESN network which consists of an input, reservoir, and readout layer. The input works as a buffer, whereas the reservoir and readout layer are dedicated to feature extraction and classification, respectively.
@@ 2. Overview of Echo state Network
The ESN is a class of reservoir networks originally proposed by Jaeger (Jaeger, 2001) in 2002 to circumvent challenges associated with training recurrent neural networks (RNNs) (vanishing and exploding gradient when using backpropagation through time (Gauthier et al., 2021)). The algorithm is inspired by the structure of the neocortex and attempts to model its dynamic and sequence learning (Zhang and Vargas, 2023). The high-level diagram of the ESN network, shown in Fig. 1, comprises three consecutive layers of $n_{u}$ input neurons, $n_{r}$ reservoir (or hidden) neurons, and $n_{o}$ readout (or output) neurons. The input layer serves as a buffer, holding the information presented to the network. The reservoir works as a feature extractor that provides non-linear dynamics and memory, enabling the network to process information of a temporal nature. The output layer is a linear classifier. When the ESN algorithm is used for time-series forecasting, it operates into three phases: initialization, prediction, and training, summarized in Algorithm-I. During the initialization phase (Line2 - Line5), the strength of the synaptic connections (weights) are randomly initialized and features associated with the structure of the network, such as the reservoir feedback connections, are set. Besides the synaptic initialization, all the non-zero elements of the reservoir weight matrix are scaled in width to get the network operating at the edge of chaos. Once the initialization phase is finished, which occurs only once, the prediction phase begins. Here, the input samples ($U=\{u^{<1>},u^{<2>},....,u^{<n_{t}>}\}$) are sequentially presented to the network where they initially get multiplied by the synaptic weights connecting the reservoir and input layer, $W_{ri}\in\mathbb{R}^{n_{r}\times n_{u}}$, and then summed up. Concurrently, within the reservoir layer, the outputs of the reservoir neurons from the previous time-step ($x^{<t-1>}$) are multiplied by their corresponding weights of feedback connections, $W_{rr}\in\mathbb{R}^{n_{r}\times n_{r}}$, and summed up. The weighted sums via both input and feedback connections are added and then subjected to a hyperbolic tangent tanh activation function ($g$) to introduce non-linearity. The output of the tanh function, $\hat{x}^{<t>}$, represents the reservoir neuron’s internal state, which is used along with the leakage rate ($\delta$) to determine the final output of the reservoir neurons, $x^{<t>}$, see [Line:8-9]. The reservoir neurons in this work capture the leaky-integrated discrete-time continuous-value feature to ensure broader control over the reservoir update dynamics. The output of the reservoir is relayed to the readout layer via the weighted synaptic connections, $W_{or}\in\mathbb{R}^{n_{o}\times n_{r}}$, and the weighted-sum is fed to a sigmoid activation function ($f$) to compute the predicted value. Once the predicted value is computed, the training phase starts to minimize the error between the predicted and targeted values (for forecasting, the targeted value is $u^{<t+n_{p}>}$ in the time-series data, where $n_{p}$ denotes the number of prediction steps). Unlike other networks, the training in the ESN is confined solely to the output layer. This makes the algorithm well-known for its fast training, thereby making it attractive to numerous applications. There are various training algorithms that can be used to train the ESN network, such as backpropagation-decorrelation, force learning (Sussillo and Abbott, 2009), ridge regression, etc., but the most common approach is the ridge regression (Lukoševičius, 2012). Ridge regression involves finding the optimal set of weights that minimize the squared error between the outputs of the network ($\hat{Y}$) and ground truth labels ($Y$). It is carried out by multiplying the Moore-Penrose generalized inverse of the reservoir output ($X\in\mathbb{R}^{n_{r}\times(n_{t}-n_{p})}$) by ground truth labels ($Y\in\mathbb{R}^{n_{o}\times(n_{t}-n_{p})}$). Orthogonal projection and Tikhonov regularization ($\lambda$) are recommended to be employed here, see (1).
@@
Input: $\vec{u}^{t}\in\mathbb{R}^{n_{u}}_{[-1,1]}$, where $\vec{u}^{t}\subset U$ and $U\in\mathbb{R}^{n_{u}\times n_{t}}_{[-1,1]};$                                /* $n_{u}$: Input features count */
Output: $\vec{\hat{y}}^{t+n_{p}}\in\mathbb{R}^{n_{c}}_{[0,1]}$;
@@
2 $W_{ri}\sim$ rand.uniform, where $W_{ri}\in\mathbb{R}^{n_{r}\times n_{u}}_{[-1,1]}$ ;
@@ 5 $count\leftarrow 0$ ;
7       // Prediction stage
8       $\hat{x}^{t}\leftarrow g(W_{ri}\odot u^{t}+W_{rr}\odot x^{t-1})$;                                                  /* g is a tanh activation function */
9       ${x}^{t}\leftarrow(1-\delta)\times x^{t-1}+\delta\times\hat{x}^{t}$ ;
10       $\hat{y}^{t}\leftarrow f(W_{or}\odot x^{t})$ ;
11       // Learning stage:
@@ 10       $\hat{y}^{t}\leftarrow f(W_{or}\odot x^{t})$ ;
13             $Er^{t}\leftarrow\hat{y}^{t}-y^{t}$;
14             $grad\leftarrow grad+x^{t}\otimes Er^{t}$;
@@ 16                   $grad[i]\leftarrow 0$ when  $grad[i]<\Theta,\forall i$ ;
17                   $W_{or}\leftarrow W_{or}-\alpha\frac{grad}{n_{up}}+\lambda W_{or}$ ;
@@ 22      $count\leftarrow count+1$ ;
While such a training approach is effective and recommended to train ESN, it is computationally costly and requires massive memory to store thousands of labels and internal states (reservoir outputs). One may resort to the recursive least squares (RLS) to mitigate the compute and memory requirements, but it is not a feasible option for edge devices with stringent resource constraints. Furthermore, most of the aforementioned training approaches suffer from a divergence problem when not trained for several time steps, generally caused by the accumulation of small errors. Such a problem is inevitable and can not be avoided by changing the structure of the network, but can be mitigated by bringing the reservoir dynamic to its original state by representing ground truth input to the reservoir. However, such an approach may work when dealing with stationary data but not real-world data of a non-stationary nature. Therefore, to overcome the challenges associated with finding the optimal set of parameters that ensure the best network performance while securing fast learning and adaptation, we use the online least mean square (LMS) algorithm due to its simplicity and minimal use for computing resources and storage. The LMS is used with weight decay regularization (LMS+L2) to optimize the ESN performance even further as it is exposed to new events, see [lines:13-20]. One may notice that the gradient is sparsified [Line:16] to avoid meaningless changes in the network parameters and to enhance the learning process in hardware. Furthermore, the weights are tuned frequently as set by the parameter $n_{up}$, rather than every iteration. This will expedite network adaptation and minimize energy consumption.
@@ 3. System Design and Implementation
The general architecture of an ESN, as covered in the overview section, is composed of consecutive layers of fully and sparsely connected neurons. In hardware, the same architecture is captured, but inter-layer connections are buffered to control data flow (see Fig. 2). During the forward pass, the input signals are sampled and held, and thereafter presented to the reservoir layer via weighted connections modeled by memristor devices. Once the weighted sum of the input reaches the reservoir neurons, collectively with weighted reservoir neuron responses from the previous time step, the reservoir neuron non-linear response is determined. The output of the reservoir is classified using the readout layer, which is typically trained in an online fashion as alluded to earlier.
@@
The input layer is equipped with a sample-and-hold (S/H) [^1] To ensure high energy efficiency and wide input range, we used the S/H circuit presented in (O’Halloran and Sarpeshkar, 2004). In the input layer, the circuit is modified to generate a differential output and to hold multiple clones of the sampled input, which will be used during the prediction and training phases. circuit to discretize the continuous time-series input signal and to hold it until the reservoir neurons are ready to receive a stimulus. The stimulus to the reservoir neurons is presented via a set of memristor devices integrated into the crossbar structure. These memristors model the synaptic weights bridging the input-reservoir layers. Using one memristor to model each synaptic weight in the network is not common, as it can solely model unipolar weights. Thus, two cross-coupled memristors (2M structure) are typically used, where the net difference between their conductance results in a bipolar weight representation. Using two memristors gives rise to double the resources utilized in each layer, leading to uncompact and less energy-efficient design. To ensure efficiency, one may resort to a 1M1R structure. In 1M1R, ($n_{u}\times n_{r}+1$) crossbar is used, where each memristor in the crossbar corresponds to one weight representation and the additional column is reserved for the reference resistor (or untunable memristor). Hence, the net difference between the conductance of any tunable memristor and the corresponding reference resistor results in bipolar weight representation (Zyarah and Kudithipudi, 2018). While the 1M1R structure uses fewer resources, it suffers from limited dynamic range and may not be favorable when implementing streaming networks due to the memristor endurance limitations. In this work, we adopted the 2M structure and attempted to leverage network and device inherent features to minimize resource usage while enhancing network robustness and performance.
@@
Figure 2. The schematic of the developed ESN accelerator, including input, reservoir, and readout layer. Each layer comprises sample-and-hold circuits to discretize and temporally hold the time-series data, memristor devices to emulate the synaptic weights, and neuron circuits (leaky-integrated discrete-time continuous-value neurons in the reservoir and point neurons in the output layer). In the output layer additional unit, training circuitry, is employed to enable in-situ learning.
(2) $\hat{x_{i}}^{<t>}$ $=g(u^{<t>}\frac{1}{M^{-}_{1}-M^{+}_{1}}+x^{<t-1>}_{1}\frac{1}{M^{-}_{2}-M^{+}_{2}}$
$\quad+x^{<t-1>}_{2}\frac{1}{M^{-}_{3}-M^{+}_{3}}+.....])$
(3) $x_{i}^{<t>}=\hat{x}^{<t>}{M_{f}}+x^{<t-1>}\frac{M_{f}}{M_{x}}$
@@
Once the stimulus reaches the reservoir layer, the updated value of each reservoir neuron’s internal state $\hat{x}^{<t>}$ is first computed as the sum of inputs ($u^{<t>}$) weighted by non-tunable memristor devices (random-fixed weights), refer to (2). Then, the non-linearity approximated by piece-wise $tanh$ function, which is conventionally achieved by leveraging the linear operating region and limited bias of pMOS-input operational amplifiers (Op-Amps), is applied (see Fig. 3-(a)). While the Op-Amps provide a virtual ground to the crossbar bit-lines, eliminating other factors impacting weight representation with memristors, using Op-Amps to approximate the non-linearity results in high power consumption and area footprint. To circumvent this issue, we exploit the virtual ground provided by the leaky integrated neuron circuit and use a diode-based limiter to approximate the non-linearity Fig. 4. This reduces the power consumption and area of the reservoir layer by 2$\times$ and 170$\times$, respectively. To capture the leaky-integrated feature, the result of the non-linearity function is multiplied by the leakage rate ($\delta$) modeled by ${M_{f}}$, and then added to reservoir neuron output from the previous time step multiplied by ($1-\delta$), which is modeled by $\frac{M_{f}}{M_{x}}$. The final output of the reservoir neuron will be relayed to the output layer and feedback circuit. In the feedback circuit, the activation of each reservoir neuron is sampled and held to be used as feedback in the upcoming iterations. However, due to the fact that the activation of the reservoir neuron is a result of tanh function, it can be either positive or negative. Storing a negative value in a S/H circuit is challenging and not feasible. To this end, we superimposed the output of each reservoir neuron on a fixed DC offset to bring all outputs to positive ranges during the sampling phase. The DC offset is canceled once the stored values are represented to the reservoir. Once the output of the reservoir layer is ready, it is multiplied by the corresponding weights, and the final output of the ESN is computed. Then the training phase starts, which will be discussed in detail in the next subsection.
Figure 3. Memristor-based reservoir neuron, in which dual-memristor devices are used to emulate bi-polar weights. The non-linearity of the neuron is approximated by: a) Op-Amps, b) diodes.
Figure 4. Approximation of the tanh activation function achieved by: i) levering the linear operating region and limited bias of pMOS-input 3-stage Op-Amp implemented in standard 65nm technology node, ii) using a diode-based limiter.

### v2

@@ Abstract.
Pushing the frontiers of time-series information processing in the ever-growing domain of edge devices with stringent resources has been impeded by the systems’ ability to process information and learn locally on the device. Local processing and learning of time-series information typically demand intensive computations and massive storage as the process involves retrieving information and tuning hundreds of parameters back in time. In this work, we developed a memristor-based echo state network accelerator that features efficient temporal data processing and in-situ online learning. The proposed design is benchmarked using various datasets involving real-world tasks, such as forecasting the load energy consumption and weather conditions. The experimental results illustrate that the hardware model experiences a marginal degradation in performance as compared to the software counterpart. This is mainly attributed to the limited precision and dynamic range of network parameters when emulated using memristor devices. The proposed system is evaluated for lifespan, robustness, and energy-delay product. It is observed that the system demonstrates reasonable robustness for device failure below 10%, which may occur due to stuck-at faults. Furthermore, 247$\times$ reduction in energy consumption is achieved when compared to a custom CMOS digital design implemented at the same technology node.
@@ 1. Introduction
Empowering edge devices with the capability to process and learn continuously locally on the device is essential when handling stationary and non-stationary time-series data, as remote cloud services have proven to be slow and unsecured (Kim et al., 2019). While such an objective enhances reliability and adaptability to edge constraints, it is challenging to satisfy in devices with limited resources. In this work, the local processing and learning from time-series data is explored within the context of reservoir networks, particularly echo state network (ESN) (Jaeger, 2001). ESN is developed to handle continuous data of stationary and non-stationary nature and has powered a large number of applications, including prosthetic finger control (Zyarah et al., 2023), stock market prediction (Wang et al., 2021), cyber-attack and anomaly detection (Hamedani et al., 2017; Ullah et al., 2022), weather forecasting (De et al., 2023), and modeling dynamic motions in bio-mimic robots (Soliman et al., 2021). The network features efficient temporal data processing and unrivaled training speed (Cerina et al., 2020; Bianchi et al., 2020). Furthermore, it does not suffer from the known training problems in most recurrent neural networks, such as vanishing and exploding gradients, which occur due to the recurrent connections and gradient backward propagation through time (Zhang and Vargas, 2023). Thus, running ESN on edge devices using dedicated custom-designed hardware not only enables the algorithm to make temporal decisions in real-time but also yields a significant improvement in power efficiency and throughput (Kudithipudi et al., 2016; Gauthier et al., 2021).
@@
These reasons urge the research community to investigate the physical implementation of ESN. There are several digital (Honda and Tamukoh, 2020; Kleyko et al., 2020; Gan et al., 2021) and mixed-signal implementations of the ESN networks in the literature using memristors devices, MOSFET switching, and photonics (Sorokina, 2020; Kume et al., 2020; Wlazlak et al., 2020; Zhang and Vargas, 2023). However, this work will place emphasis on the memristor-based implementations, owing to their numerous advantages over other approaches, such as the integration with CMOS devices and fabrication using the standard foundry process (Jiang et al., 2019). Furthermore, the memristor intrinsic non-linearity feature enhances the reservoir separability and echo state properties (Moon et al., 2019). A recent attempt at memristive ESN, including a hardware implementation of the ESN network using a memristor double-crossbar array, was presented by Hassan et al. in 2017.. The proposed design is evaluated for autonomous signal generation using the Mackey-Glass dataset and the simulation is done in MATLAB. In 2019, Wen et al. presented a memristor-based ESN trained in online fashion using least mean square (Wen et al., 2018). The network is evaluated for short-term power load forecasting. In the same year, a memristor-based reservoir computing system, which leverages the concept of the virtual node, was introduced by Moon at al. (Moon et al., 2019). The virtual node concept involves building a single reservoir node and using it to virtually emulate a chain of other nodes in the reservoir. The proposed system features high-power efficiency but suffers from high latency and short lifespan, and it is demonstrated in speech recognition and time-series forecasting. In 2022, the same concept was adopted to build a cyclic reservoir computing system based on memristor devices. The proposed design is implemented in a 65nm technology node and verified on hand-written vowel recognition task (Liang et al., 2022). Lately, Nair et al. proposed a memristive ESN with an extended synaptic sampling machine (SSM) to ensure synaptic stochasticity and power efficiency (Nair et al., 2023). The proposed neural system is benchmarked for classification tasks using ECG, MNIST, Fashion MNIST, and CIFAR10.
@@
While a few of the ESN implementations are equipped with in-situ learning to endow the ESN network with fast learning and adaptation capabilities, none extend the learning to cover the structural plasticity (synaptogenesis), which is imperative for controlling the network’s sparsity level. To the best of our knowledge, few of them utilize leaky-integrated discrete-time continuous-value neurons in the reservoir layer. Having leaky-integrated neurons is essential when dealing with temporal data as it controls the speed of the reservoir update dynamics and eventually the duration of the short-term memory of ESN (Lukoševičius, 2012). Besides the lack of using leaky-integrated neurons, none of the aforementioned implementations take into consideration the limited precision of signal digitization and undesired leakage of the sample and hold circuits during network training and testing. This paper addresses these problems with key contributions summarized as follows:
@@
The rest of the paper is organized as follows: Section II discusses the theory of ESN network and training procedure. Section III presents the system design and implementation of the ESN. The design methodology is introduced in Section IV. Sections V and VI discuss the results and conclude the paper, respectively.
Figure 1. A high-level diagram of the ESN network which consists of an input, reservoir, and readout layer. The input works as a buffer, whereas the reservoir and readout layers are dedicated to feature extraction and classification, respectively.
@@ 2. Overview of Echo state Network
The ESN is a class of reservoir networks originally proposed by Jaeger (Jaeger, 2001) in 2002 to circumvent challenges associated with training recurrent neural networks (RNNs) (vanishing and exploding gradient when using backpropagation through time (Gauthier et al., 2021)). The algorithm is inspired by the structure of the neocortex and attempts to model its dynamics and sequence learning (Zhang and Vargas, 2023). The high-level diagram of the ESN, shown in Fig. 1, comprises three consecutive layers of $n_{u}$ input neurons, $n_{r}$ reservoir (or hidden) neurons, and $n_{o}$ readout (or output) neurons. The input layer serves as a buffer, holding the information presented to the network. The reservoir works as a feature extractor that provides non-linear dynamics and memory, enabling the network to process information of a temporal nature. The output layer is a linear classifier. When the ESN algorithm is used for time-series forecasting, it operates in three phases: initialization, prediction, and training, summarized in Algorithm-I. During the initialization phase (Lines 2-5), the strength of the synaptic connections (weights) is randomly initialized and the features associated with the structure of the network, such as the reservoir feedback connections, are set. Besides the synaptic initialization, all non-zero elements of the reservoir weight matrix are scaled in width to get the network operating at the edge of chaos. Once the initialization phase is finished, which occurs only once, the prediction phase begins. Here, the input examples ($U=\{\vec{u}^{1},\vec{u}^{2},....,\vec{u}^{n_{m}}\}$) are presented to the network in an online fashion. Given an example, $\vec{u}\in\mathbb{R}^{n_{u}\times n_{t}}$, its features are sequentially presented to the network where they initially get multiplied by the synaptic weights connecting the reservoir and input layer, $W_{ri}\in\mathbb{R}^{n_{r}\times n_{u}}$, and then summed up. Concurrently, within the reservoir layer, the outputs of the reservoir neurons from the previous time step ($x^{<t-1>}$) are multiplied by their corresponding feedback connection weights, $W_{rr}\in\mathbb{R}^{n_{r}\times n_{r}}$, and summed up. The weighted sums from both input and feedback connections are added and then subjected to a hyperbolic tangent tanh activation function ($g$) to introduce non-linearity. The output of the tanh function, $\hat{x}^{<t>}$, represents the reservoir neuron’s internal state, which is used along with the leakage rate ($\delta$) to determine the final output of the reservoir neurons, $x^{<t>}$, see [Lines 8-9]. The reservoir neurons in this work capture the leaky-integrated discrete-time continuous-value feature to ensure broader control over the reservoir update dynamics. The outputs of the reservoir neurons are then transmitted to the readout layer via the weighted synaptic connections, $W_{or}\in\mathbb{R}^{n_{o}\times n_{r}}$, and the weighted sum is subsequently fed to a sigmoid activation function ($f$) to compute the predicted value. Once the predicted value is computed, the training phase starts to minimize the error between the predicted and targeted values (for forecasting, the targeted value is $u^{<t+n_{p}>}$ in the time-series data, where $n_{p}$ denotes the number of prediction steps). Unlike other networks, the training in the ESN is confined solely to the output layer. This makes the algorithm well-known for its fast training, thereby making it attractive to numerous applications. Various training algorithms can be used to train the ESN, such as backpropagation-decorrelation, force learning (Sussillo and Abbott, 2009), ridge regression, etc., but the most common approach is ridge regression (Lukoševičius, 2012). Ridge regression involves finding the optimal set of weights that minimize the squared error between the outputs of the network ($\hat{Y}$) and ground truth labels ($Y$). It is carried out by multiplying the Moore-Penrose generalized inverse of the reservoir output ($X$) by ground truth labels. Orthogonal projection and Tikhonov regularization ($\lambda$) are recommended to be employed here to overcome overfitting and enhance stability, see (1).
@@
Input: $\vec{u}\in\mathbb{R}^{n_{u}\times n_{t}}_{[-1,1]}$, where $\vec{u}\subset U$ and $U\in\mathbb{R}^{n_{u}\times n_{m}\times n_{t}}_{[-1,1]};$                                /* $n_{m}$: Number of examples */
Output: $\vec{\hat{y}}\in\mathbb{R}^{n_{c}\times(n_{t}-n_{p})}_{[0,1]}$;
@@
2 $W_{ri}\sim$ rand.uniform, where $W_{ri}\in\mathbb{R}^{n_{r}\times n_{u}}_{[-1,1]}$ ;                                                /* $n_{u}$: Input features count */
@@ 5 $count\leftarrow 1$ ;
7       // Prediction phase
8       $\hat{x}^{<t>}\leftarrow g(W_{ri}\odot u^{<t>}+W_{rr}\odot x^{<t-1>})$;                                         /* $g$ is a tanh activation function */
9       ${x}^{<t>}\leftarrow(1-\delta)\leavevmode\nobreak\ x^{<t-1>}+\delta\leavevmode\nobreak\ \hat{x}^{<t>}$ ;
10       $\hat{y}^{<t>}\leftarrow f(W_{or}\odot x^{<t>})$ ;                                                             /* $f$ is a sigmoid activation function */
11       // Learning phase:
@@
13             $Er^{<t>}\leftarrow\hat{y}^{<t>}-y^{<t>}$;
14             $grad\leftarrow grad+x^{<t>}\otimes Er^{<t>}$;
@@ 16                   $grad[i]\leftarrow 0$ when  $grad[i]<\Theta,\forall i$ ;
17                   $W_{or}\leftarrow W_{or}-\alpha\leavevmode\nobreak\ \frac{grad}{n_{up}}+\lambda\leavevmode\nobreak\ W_{or}$ ;
@@ 22      $count\leftarrow count+1$ ;
While such a training approach is effective and recommended for training ESN, it is computationally costly and requires massive memory to store thousands of labels and states (reservoir outputs). One may resort to the recursive least squares (RLS) to mitigate the computational and memory requirements, but this approach is not feasible for edge devices with stringent resource constraints. It is important to highlight that the aforementioned training approaches suffer from a divergence problem when not trained for several time steps, generally caused by the accumulation of small errors. Such a problem is inevitable and cannot be avoided by altering the structure of the network, but can be mitigated by bringing the reservoir dynamic to its original state by representing ground truth input to the reservoir (Moon et al., 2019). However, such an approach may work when dealing with stationary data but not real-world data of a non-stationary nature. Therefore, to overcome the challenges associated with finding the optimal set of parameters that ensure the best network performance while securing fast learning and adaptation, we use the online least mean square (LMS) algorithm due to its simplicity and minimal use for computational resources and storage. The LMS is used with weight decay regularization (LMS+L2) to optimize ESN performance even further as it is exposed to new events, see [lines 13-20]. Besides regularization, the gradients are sparsified [Line 16] to avoid insignificant changes in the network parameters and to enhance the learning process in hardware. Furthermore, the weights are tuned frequently according to the parameter $n_{up}$, rather than every iteration. This will expedite network adaptation and minimize energy consumption.
@@ 3. System Design and Implementation
The general architecture of an ESN, as covered in the overview section, is composed of consecutive layers of fully and sparsely connected neurons. In hardware, the same architecture is captured, but inter-layer connections are buffered to control data flow (see Fig. 2). During the forward pass, the input signals are sampled and held, and then presented to the reservoir layer via weighted connections modeled by memristor devices. Once the weighted sum of the input reaches the reservoir neurons, collectively with weighted reservoir neuron responses from the previous time step, the reservoir neuron non-linear response is determined. The output of the reservoir is classified using the readout layer, which is trained in an online fashion as alluded to earlier.
@@
The input layer is equipped with a sample-and-hold (S/H) circuit [^1] To ensure high energy efficiency and a wide input range, we used the S/H circuit presented in (O’Halloran and Sarpeshkar, 2004). The circuit is modified to generate a differential output and to hold multiple clones of the sampled input, which will be used during the prediction and training phases. to discretize the continuous time-series input signal and to hold it until the reservoir neurons are ready to receive a stimulus. The stimulus to the reservoir neurons is presented via a set of memristor devices integrated into the crossbar structure. These memristors model the synaptic weights bridging the input-reservoir layers. Using one memristor to model each synaptic weight in the network is not common, as it can solely model unipolar weights. Thus, two cross-coupled memristors (2M structure) are typically used, where the net difference between their conductance results in a bipolar weight representation. Using two memristors doubles the resources utilized in each layer, leading to uncompact and a less energy-efficient design. To ensure efficiency, one may resort to a 1M1R structure. In 1M1R, ($n_{u}\times n_{r}+1$) crossbar is used, where each memristor in the crossbar corresponds to one weight representation and the additional column is reserved for the reference resistor (or untunable memristor). Hence, the net difference between the conductance of any tunable memristor and the corresponding reference resistor results in bipolar weight representation (Zyarah and Kudithipudi, 2018). While the 1M1R structure uses fewer resources, it suffers from a limited dynamic range and may not be favorable when implementing streaming networks due to memristor endurance limitations. In this work, we adopted the 2M structure and attempted to leverage network and device inherent features to minimize resource usage while enhancing network robustness and performance.
@@
Figure 2. The schematic of the developed ESN accelerator, including input, reservoir, and readout layers. Each layer comprises sample-and-hold circuits to discretize and temporally hold the time-series data, memristor devices to emulate the synaptic weights, and neuron circuits (leaky-integrated discrete-time continuous-value neurons in the reservoir and point neurons in the output layer). In the output layer, an additional unit, the training circuitry, is employed to enable in-situ learning.
(2) $\hat{x_{i}}^{<t>}=g[(\frac{R_{f}}{M^{-}_{i1}}-\frac{R_{f}}{M^{+}_{i1}})\leavevmode\nobreak\ u^{<t>}+(\frac{R_{f}}{M^{-}_{i2}}-\frac{R_{f}}{M^{+}_{i2}})\leavevmode\nobreak\ x^{<t-1>}_{1}+....+(\frac{R_{f}}{M^{-}_{i(n_{h}+1)}}-\frac{R_{f}}{M^{+}_{i(n_{h}+1)}})\leavevmode\nobreak\ x^{<t-1>}_{n_{h}}]$
(3) $x_{i}^{<t>}\simeq\underbrace{\frac{M_{z}||M_{y}}{(M_{z}||M_{y})+M_{x}}}_{\delta}\hat{x}^{<t>}+\underbrace{\frac{M_{z}||M_{x}}{(M_{z}||M_{x})+M_{y}}}_{1-\delta}x^{<t-1>}$
@@
When the stimulus reaches the reservoir layer, the internal states of all the reservoir neurons, $\hat{x}^{<t>}$, are updated. This process begins by computing the sum of inputs ($u^{<t>}$ and $x^{<t-1>}$) weighted by non-tunable memristor devices (random-fixed weights), as described in (2). Then, the non-linearity, represented by $tanh$ function, is approximated by leveraging the linear operating region and limited bias of pMOS-input operational amplifiers (Op-Amps). After determining the internal states of the reservoir neuron, the leaky-integrated feature is captured to assess how much information from the previous activation should be carried over to the current hidden state. Implementing this feature in hardware may involve using at least one additional Op-Amp, leading to high power consumption and a large footprint area. To circumvent this issue, we introduce the so called leakage cell. The leakage cell combines the internal state and the previous activations using three memristors connected as shown in Fig. 2. These memristors model the leakage terms, with the internal state being multiplied by the leakage rate ($\delta$), modeled by $\frac{M_{z}||M_{y}}{(M_{z}||M_{y})+M_{x}}$, and the previous reservoir neuron output being multiplied by ($1-\delta$), modeled by $\frac{M_{z}||M_{x}}{(M_{z}||M_{x})+M_{y}}$ [^2] The resistance of the $M_{z}$ memristor should be set to be really large as compared to $M_{x}$ and $M_{y}$ so that the approximated $\delta$ and $1-\delta$ always sum up to one.. This approach reduces the power consumption and area of the reservoir layer by more than 2$\times$ and 170$\times$, respectively, compared to using Op-Amps. However, once the final outputs of the reservoir neurons are ready, they will be relayed to the feedback circuit. In the feedback circuit, the activation of each reservoir neuron is bypassed to the output layer to compute the final network output, $\hat{y}$, and sampled and held to be used as feedback in the upcoming iterations. Due to the fact that the activation of the reservoir neuron is a result of tanh function, it can be either positive or negative. Storing a negative value in a S/H circuit is challenging and not feasible. To this end, we superimposed the output of each reservoir neuron on a fixed DC offset to bring all outputs to positive ranges during the sampling phase. The DC offset is canceled once the stored values are represented to the reservoir. Once the output of the reservoir layer is ready, it is multiplied by the corresponding weights, and the final output of the ESN is computed. Then the training phase starts, which will be discussed in detail in the next subsection.